Calculate the de Broglie wavelength of a non-relativistic proton with a kinetic energy of 10 eV.
1
Expert's answer
2021-10-31T18:12:19-0400

Given:

"E_k=10\:\rm eV=1.6*10^{-18}\: J"


The de Broglie wavelength is given by

"\lambda=\frac{h}{p}=\frac{h}{\sqrt{2mE_k}}""\lambda=\frac{6.67*10^{-34}}{\sqrt{2*1.67*10^{-27}*1.6*10^{-18}}}=9.1*10^{-12}\:\rm m"